The number of elements in the Power set P(S) of the set S = { { Φ} , 1, { 2, 3 }} is?







1
Expert's answer
2021-08-11T08:55:29-0400

Answer: 8.The power set of any set S is the set of all subsets of S, including the empty set and the set S itself. The number of elements will be 2n where n is the number of elements of S. The power set is denoted as P(S). Let the number of elements of the power set of S be x, then
x"=2^n=2^3=8"